# Large Language Models as Urban Residents: An LLM Agent Framework for Personal Mobility Generation

### v1

@@ Abstract
This paper introduces a novel approach using Large Language Models (LLMs) integrated into an agent framework for flexible and efficient personal mobility generation. LLMs overcome the limitations of previous models by efficiently processing semantic data and offering versatility in modeling various tasks. Our approach addresses the critical need to align LLMs with real-world urban mobility data, focusing on three research questions: aligning LLMs with rich activity data, developing reliable activity generation strategies, and exploring LLM applications in urban mobility. The key technical contribution is a novel LLM agent framework that accounts for individual activity patterns and motivations, including a self-consistency approach to align LLMs with real-world activity data and a retrieval-augmented strategy for interpretable activity generation. In experimental studies, comprehensive validation is performed using real-world data. This research marks the pioneering work of designing an LLM agent framework for activity generation based on real-world human activity data, offering a promising tool for urban mobility analysis.
@@ 1 Introduction
Under the 2030 Sustainable Development Agenda, the United Nations emphasizes the crucial role of sustainable cities and human settlements. The study of personal mobility like individual activity trajectory can provide valuable semantic insights into urban mobility and support this objective. Specifically, modeling personal mobility opens up numerous applications for building a sustainable community, including proactive traffic management and the design of comprehensive urban development strategies Batty et al. (2012); Batty (2013); Zheng (2015). In particular, generating reliable activity trajectories can be a promising and efficient way to exploit individual activity data. On one hand, learning to generate activity trajectory leads to a thorough understanding of activity patterns, enabling flexible simulation of urban mobility. On the other hand, while individual activity trajectory data is abundant due to advances in telecommunications, its practical use is often limited by privacy concerns. In this sense, generated data can provide a viable alternative that offers a balance between utility and privacy.
@@
While advanced data-driven learning-based method offer various solutions to generate synthetic individual trajectoriesHuang et al. (2019); Zhang et al. (2020); Feng et al. (2020); Choi et al. (2021), the generated data only imitates real-world data from the data distribution perspective rather than semantics, rendering them less effective in simulating or interpretation activities in novel or unforeseen scenarios with a significantly different distribution (e.g., a pandemic). Thus, in this study, to explore a more intelligent and efficient activity generation, we propose to establish an agent framework by exploiting the emerging intelligence of large language models (LLMs), as illustrated by Fig. 1. LLMs have gained widespread usage in furthering our understanding of humans and society in a multitude of disciplines, such as political science Argyle et al. (2023) and economy Aher et al. (2022), and thus have been employed as agents in a variety of social simulations Xi et al. (2023); Wu et al. (2023); Gao et al. (2023). More specifically, LLMs present two significant advantages over previous models when applied to activity trajectory generation:
@@
  • Semantic Interpretability. Unlike previous models, which have predominantly depended on structured data (e.g., GPS coordinates-based trajectory data) for both calibration and simulation Jiang et al. (2016); Pappalardo and Simini (2018), LLMs exhibit proficiency in interpreting semantic data (e.g., activity trajectory data). This advantage significantly broadens the scope for incorporating a diverse array of data sources into generation processes, thereby enhancing the models’ ability to understand and interact with complex, real-world scenarios in a more nuanced and effective manner.
@@
  • Model Versatility. Although other data-driven methods manage to learn such dynamic activity patterns for generation, their capacity is limited for generation under unseen scenarios. On the contrary, LLMs have shown remarkable versatility in dealing with unseen tasks, especially the ability to reason and decide based on available information OpenAI (2022). This competence enables LLMs to offer a diverse and rational array of choices, making it a promising and flexible approach for modeling personal mobility patterns.
@@
Despite these benefits, ensuring that LLM aligns effectively with real-world situations continues to be a significant challenge Xi et al. (2023). This alignment is particularly crucial in the context of urban mobility, where the precision and dependability of LLM outputs are essential for the efficacy of any urban management derived from them. In this study, our aim is to address this challenge by investigating the following research questions: RQ 1: How can LLM be effectively aligned with semantically rich data about daily individual activities? RQ 2: What are the effective strategies for achieving reliable and meaningful activity generation using LLM? RQ 3: What are the potential applications of LLM in enhancing urban mobility?
@@
Our methodology revolves around two phases: (1) self-consistent activity pattern identification and (2) motivation-driven activity generation. In Phase 1, we extract and evaluate activity patterns from historical data and semantic analysis, using LLM agents to generate diverse and personalized activity scenarios. This process involves identifying habitual activity patterns, which are key factors influencing individual mobility. In Phase 2, we develop two interpretable retrieval-augmented strategies, along with the patterns identified in Phase 1, to instruct LLM agents to work out summarized motivations, such as evolving interests or situational needs. The patterns and motivations faciliate the agents to generate daily activity trajectories. The generation benefits from the LLM’s capability to process extensive datasets and semantic richness, allowing for a nuanced simulation of personal mobility with considerations for habitual behaviors, situational motivations, and temporal dynamics.
@@
We evaluate the proposed framework using GPT-3.5 APIs over a personal activity trajectory dataset of Tokyo. The results demonstrate the capability of our framework to align LLM agents with semantically rich data for generating individual daily activities, highlighting that our framework excels in temporal aspects of personal mobility generation, with spatial-temporal metrics showing its strength in capturing the essence of daily routines over mere location accuracy. The comparison with baselines like Markov models, LSTM, and DeepMove Feng et al. (2018), underscores the advanced generative performance of our framework. Moreover, the application of the framework in simulating urban mobility under specific contexts, such as the pandemic scenario, reveals its potential to adapt to external factors and generate realistic activity patterns.
@@
To the best of our knowledge, this study is one of the pioneering works in developing an LLM agent framework for generating activity trajectory based on real-world data. We summarize our contributions as follows: (1) We introduce a novel LLM agent framework for urban mobility generation featuring semantic richness. (2) Our framework introduces a self-consistency evaluation to ensure that the output of the LLM aligns closely with real-world data on daily activities. (3) To generate daily activity trajectories, our framework integrates activity patterns with summarized motivations, with two interpretable retrieval-augmented strategies aimed at producing reliable activity trajectories. (4) By using real-world personal activity data, we validate the effectiveness of our framework and explore its utility in urban mobility analysis.
@@ 2.1 Personal Mobility Generation
Activity trajectory generation offers a valuable perspective for understanding urban mobility. Benefiting from the advanced communication technologies, recent research can explore urban mobility from massive geotagged data. Based on vast call detailed records, Jiang et al.Jiang et al. (2016) built a mechanistic modeling framework to generate individual activities in high spatial-temporal resolutions, but it is limited to three activity types: work, home, and others. Pappalardo and SiminiPappalardo and Simini (2018) enhanced the Exploration and Preferential Return (EPR) model for activity generation, which employs Markov modeling to estimate the probability of individuals visiting specific locations. Besides, deep learning has become a robust tool for modeling the complex dynamics of traffic  Liu et al. (2018); Huang et al. (2019); Zhang et al. (2020); Feng et al. (2020); Choi et al. (2021); Luca et al. (2021). The primary challenge involves overcoming data-related obstacles such as randomness, sparsity, and irregular patterns Feng et al. (2018); Yuan et al. (2023, 2022); Long et al. (2023). For example, Feng et al.Feng et al. (2018) proposed attentional recurrent networks to handle personal preference and transition regularities. Yuan et al.Yuan et al. (2022) leveraged deep learning combined with neural differential equations to address the challenges of randomness and sparsity inherent in irregularly sampled activities for activity trajectory generation.
Figure 2: Illustration of LLMob, the proposed LLM agent framework for personal Mobility generation.
2.2 LLM Agents
How to take advantage of LLMs for real-world tasks has become an attractive research topic since the LLMs present human-like intelligence. While some studies Wei et al. (2022); Wang et al. (2022) focus on exploiting their linguistic capabilities, exploring how to treat LLM as autonomous agents in specific scenarios can lead to more diverse and promising applications Wang et al. (2023); Xi et al. (2023); Wu et al. (2023); Gao et al. (2023), hence to address the issues of reliance on parameter settings and lack of adaptability in traditional agent-based simulations using rules or reinforcement learning Grimm et al. (2005); Henrich et al. (2010); Ale Ebrahim Dehkordi et al. (2023). For instance, Park et al.Park et al. (2023) first established an LLM-powered agent framework to simulate human behavior in an interactive scenario. This innovative approach demonstrates the potential of LLM to model complex social interactions and decision-making processes. Extending beyond the realm of social sciences, Mao et al.Mao et al. (2023) have adeptly utilized LLM to generate driving trajectories in motion planning tasks. In the field of natural sciences, Williams et al.Williams et al. (2023) have innovatively integrated LLM with epidemic models to simulate the spread of diseases. Furthermore, in financial research, the application of LLM in macroeconomic modeling has been explored, providing new insights into financial markets and economies Li et al. (2023). These varied applications highlight the versatility and potential of LLM to understand and model various real-world dynamics.
3 Problem Description
We study the generation of individual daily activity trajectories, each representing an individual’s activities for the whole day. We focus on the urban context, where the activities of each individual are represented as a time-ordered sequence of location choices (e.g., location-based activity) Luca et al. (2021). This sequence, $\{(l_{0},t_{0}),(l_{1},t_{1}),\ldots,(l_{n},t_{n})\}$, represents the activity trajectory of an individual (e.g., individual POI trajectories), where $(l,t)$ denotes the individual’s location $l$ at time $t$.
@@ 4 Methodology
By modeling individuals within an urban environment as LLM agents, we present LLMob, an LLM Agent Framework for Personal Mobility Generation, as illustrated in Fig. 2. LLMob is based on the assumption that an individual’s activities are primarily influenced by two principal factors: habitual activity patterns and current motivations. Habitual activity patterns, which reflect typical movement behaviors and preferences, signifying regular travel and location choices Song et al. (2010). Motivations pertain to dynamic and situational factors that influence an individual’s decisions at any given moment, such as specific needs or external influences during a particular time frame. This consideration is vital for capturing and forecasting short-term shifts in mobility patterns Aher et al. (2022). Moreover, by formulating prompts that assume specific events of concern, this framework allows us to observe the agents’ activities and responses in a variety of situations.
To construct a pipeline for activity trajectory generation, we design LLM agents with action, memory, and planning, in line with Weng (2023); Wang et al. (2023) except that external tools are not involved. Action specifies how an agent interacts with the environment and makes decisions. In LLMob, the environment contains the information collected from real-world data, and the agents act by identifying habitual activity patterns and generating trajectories. Memory includes past actions that need to be prompted to the LLM to invoke the next action. In LLMob, memory refers to the active patterns output by the agents. Planning formulates a plan or reflects over past actions to handle complex tasks. In LLMob, planning refers to the generation of summarized motivations, which are used to facilitate trajectory generation. Furthermore, we specify personas to the agents to simulate the diversity of real-world individuals Salewski et al. (2023).
4.1 Self-Consistent Activity Pattern Identification
Phase 1 of LLMob focuses on identifying activity patterns from historical data. We leverage the extracted activity patterns as essential prior knowledge for the generation of daily activities. To effectively identify activity patterns from historical data, we introduce the following two steps.
@@ 4.1.1 Pattern Extraction from Semantics and Historical Data
The first step will use LLM to derive activity patterns based on activity trajectory data (e.g. individual check-in data). As illustrated in the left panel of Fig. 2, this scheme consists of three aspects: For each person, we start by utilizing the LLM to create a variety of personas, providing the inspiring foundation for subsequent activity pattern generation. This approach also encourages the diversity of the generated activity patterns, as each persona acts as a unique prior for the generation process (e.g., Noulas et al.Noulas et al. (2011) demonstrated the significance of user clustering from activity trajectory data in producing meaningful distinctions). Meanwhile, we perform data preprocessing to extract key information from the extensive historical data. This involves identifying usual commuting distances, pinpointing typical start and end times and locations of daily trips, and concluding the most frequently visited locations of the person. It is important to note that these pieces of information are widely recognized as critical features in mobility analysis Jiang et al. (2016). Following this, both semantic elements with historical data are combined in the prompts, requiring the LLM to summarize the activity patterns for this person. By doing this, we set up a streamline to effectively bridge the gap between semantic persona characteristics and concrete historical activity trajectory data, which allows for a more personalized and interpretable representation of individual activities at one day.
@@ 4.1.2 Pattern Evaluation with Self-Consistency
As introduced before, for each person, multiple candidate patterns are generated according to the persona and historical data. Following this, the subsequent step involves assessing the consistency of these generated activity patterns to identify the most plausible one. We implement a scoring mechanism to gauge the alignment of these patterns with historical data. To achieve this objective, we define a scoring function for each candidate pattern $cp$ within the set $\mathcal{CP}$. This function evaluates $cp$ against two distinct sets of activity trajectories: the specific activity trajectories $\mathcal{T}_{i}$ of a targeted user $i$ and the sampled activity trajectories from other users $\mathcal{T}_{\sim i}$:
$score_{cp}=\sum_{t\in\mathcal{T}_{i}}r_{t}-\sum_{t^{\prime}\in\mathcal{T}_{\sim i}}r_{t^{\prime}},$ (1)
where we design an evaluation prompt to ask the LLM to generate rating scores $r_{t}$ and $r_{t^{\prime}}$. Specifically, the LLM is prompted to assess the degree of preference for a given trajectory based on the candidate pattern. Ideally, the LLM should assign a higher $r_{t}$ for data from the targeted user and a lower $r_{t^{\prime}}$ for data from other users. This scheme requires self-consistency in the pattern: the activity pattern derived from the activity trajectory data should be consistent with the data during the evaluation. The algorithm for this step is summarized as shown in Alg. 1.
Input : Number of personas $C$, activity trajectories $\mathcal{T}$
Output : Best activity pattern $best\_pattern$
1 Initialize: Empty candidate pattern set $\mathcal{CP}$;
3 for $i=1$ to $C$ do
4       Formulate $p_{1}$ with prior information and persona $i$;
5       $cp\leftarrow\text{LLM}(p_{1})$;
6       $\mathcal{CP}\leftarrow\mathcal{CP}\cup\{\,cp\,\}$;
7
8Initialize a score dictionary $\mathcal{S}$ to store pattern scores;
9 foreach candidate pattern $cp$ in $\mathcal{CP}$ do
10       Initialize $score_{cp}=0$;
11       foreach activity trajectory $t$ in $\mathcal{T}$ do
12             Formulate evaluation prompt $p_{2}$ for $cp$ and $t$;
13             $score_{cp}\leftarrow score_{cp}+\text{LLM}(p_{2})$;
14      $\mathcal{S}[cp]\leftarrow score_{cp}$;
15
[…]
return $best\_pattern$
Algorithm 1 Self-Consistent Activity Pattern Identification

### v2

# font=, innertopmargin=2pt, innerbottommargin=2pt, innerleftmargin=8pt, innerrightmargin=8pt, frametitleaboveskip=2pt, frametitlebelowskip=1pt, frametitlealignment=, middlelinecolor=blue

@@ Abstract
This paper introduces a novel approach using Large Language Models (LLMs) integrated into an agent framework for flexible and effective personal mobility generation. LLMs overcome the limitations of previous models by effectively processing semantic data and offering versatility in modeling various tasks. Our approach addresses three research questions: aligning LLMs with real-world urban mobility data, developing reliable activity generation strategies, and exploring LLM applications in urban mobility. The key technical contribution is a novel LLM agent framework that accounts for individual activity patterns and motivations, including a self-consistency approach to align LLMs with real-world activity data and a retrieval-augmented strategy for interpretable activity generation. We evaluate our LLM agent framework and compare it with state-of-the-art personal mobility generation approaches, demonstrating the effectiveness of our approach and its potential applications in urban mobility. Overall, this study marks the pioneering work of designing an LLM agent framework for activity generation based on real-world human activity data, offering a promising tool for urban mobility analysis.
Source codes are available at https://github.com/Wangjw6/LLMob/ .
@@ 1 Introduction
The prevalence of large language models (LLMs) has facilitated a variety of applications extending beyond the domain of NLP. Notably, LLMs have gained widespread usage in furthering our understanding of humans and society in a multitude of disciplines, such as economy [1] and political science [2], and have been employed as agents in various social science studies [34, 33, 10]. In this paper, we target the utilization of LLM agents for the study of personal mobility data. Modeling personal mobility opens up numerous opportunities for building a sustainable community, including proactive traffic management and the design of comprehensive urban development strategies [4, 3, 40]. In particular, generating reliable activity trajectories has become a promising and effective way to exploit individual activity data [13, 6]. On one hand, learning to generate activity trajectory leads to a thorough understanding of activity patterns, enabling the flexible simulation of urban mobility. On the other hand, while individual activity trajectory data is abundant thanks to advances in telecommunications, its practical use is often limited due to privacy concerns. In this sense, generated data can provide a viable alternative that offers a balance between utility and privacy.
@@
While advanced data-driven learning-based methods offer various solutions to generate synthetic individual trajectories [13, 39, 9, 6], the generated data only imitates real-world data from the data distribution perspective rather than semantics, rendering them less effective in simulating or interpreting activities in novel or unforeseen scenarios with a significantly different distribution (e.g., a pandemic). Thus, in this study, to explore a more intelligent and effective activity generation, we propose to establish a trajectory generation framework by exploiting the emerging intelligence of LLM agents, as illustrated in Figure 1. LLMs present two significant advantages over previous models when applied to activity trajectory generation:
@@
  • Semantic Interpretability. Unlike previous models, which have predominantly depended on structured data (e.g., GPS coordinates-based trajectory data) for both calibration and simulation [15, 26], LLMs exhibit proficiency in interpreting semantic data (e.g., activity trajectory data). This advantage significantly broadens the scope for incorporating a diverse array of data sources into generation processes, thereby enhancing the models’ ability to understand and interact with complex, real-world scenarios in a more nuanced and effective manner.
@@
  • Model Versatility. Although other data-driven methods manage to learn such dynamic activity patterns for generation, their capacity is limited for generation under unseen scenarios. On the contrary, LLMs have shown remarkable versatility in dealing with unseen tasks, especially the ability to reason and decide based on available information [25]. This competence enables LLMs to offer a diverse and rational array of choices, making it a promising and flexible approach for modeling personal mobility patterns.
@@
Despite these benefits, ensuring that LLMs align effectively with real-world situations continues to be a significant challenge [34]. This alignment is particularly crucial in the context of urban mobility, where the precision and dependability of LLM outputs are essential for the efficacy of any urban management derived from them. In this study, our aim is to address this challenge by investigating the following research questions: RQ 1: How can LLMs be effectively aligned with semantically rich data about daily individual activities? RQ 2: What are the effective strategies for achieving reliable and meaningful activity generation using LLM agents? RQ 3: What are the potential applications of LLM agents in enhancing urban mobility analysis?
@@
To this end, our study employs LLM agents to infer activity patterns and motivation for personal activity generation tasks. While previous researches advocate habitual activity patterns and motivations as two critical elements for activity generation [15, 38], our proposed framework introduces a more interpretable and effective solution. By leveraging the capabilities of LLMs to process semantically rich datasets (e.g., personal check-in data), we enable a nuanced and interpretable simulation of personal mobility. Our methodology revolves around two phases: (1) activity pattern identification and (2) motivation-driven activity generation. In Phase 1, we leverage the semantic awareness of LLM agents to extract and identify self-consistent, personalized habitual activity patterns from historical data. In Phase 2, we develop two interpretable, retrieval-augmented strategies that utilize the patterns identified in Phase 1. These strategies guide LLM agents to infer underlying daily motivations, such as evolving interests or situational needs. Finally, we instruct LLM agents to play urban residents according to the obtained patterns and motivations and generate their daily activities.
@@
We evaluate the proposed framework using GPT-3.5 APIs over a personal activity trajectory dataset of Tokyo. The results demonstrate the capability of our framework to align LLM agents with semantically rich data for generating individual daily activities. The comparison with baselines, such as attention-based methods [8, 21], adversarial learning methods [6, 37], and a diffusion model [41], underscores the advanced generative performance of our framework. The observation also suggests that our framework excels in reproducing temporal and spatio-temporal aspects of personal mobility generation and interpretable activity routines. Moreover, the application of the framework in simulating urban mobility under specific contexts, such as a pandemic scenario, reveals its potential to adapt to external factors and generate realistic activity patterns.
@@
To the best of our knowledge, this study is one of the pioneering works in developing an LLM agent framework for generating activity trajectory based on real-world data. We summarize our contributions as follows: (1) We introduce a novel LLM agent framework for personal mobility generation featuring semantic richness. (2) Our framework introduces a self-consistency evaluation to ensure that the output of LLM agents aligns closely with real-world data on daily activities. (3) To generate daily activity trajectories, our framework integrates activity patterns with summarized motivations, with two interpretable retrieval-augmented strategies aimed at producing reliable activity trajectories. (4) By using real-world personal activity data, we validate the effectiveness of our framework and explore its utility in urban mobility analysis.
@@ 2.1 Personal Mobility Generation
Activity trajectory generation offers a valuable perspective for understanding personal mobility. Based on vast call detailed records, Jiang et al. built a mechanistic modeling framework to generate individual activities in high spatial-temporal resolutions. Pappalardo and Simini employed Markov modeling to estimate the probability of individuals visiting specific locations. Besides, deep learning has become a robust tool for modeling the complex dynamics of traffic  [18, 13, 39, 9, 6, 20]. The primary challenge involves overcoming data-related obstacles such as randomness, sparsity, and irregular patterns [8, 38, 37, 19]. For example, Feng et al. proposed attentional recurrent networks to handle personal preference and transition regularities. Yuan et al. leveraged deep learning combined with neural differential equations to address the challenges of randomness and sparsity inherent in irregularly sampled activities for activity trajectory generation. Recently,  Zhu et al. proposed to utilize a diffusion model to generate GPS trajectories.
2.2 LLM Agents in Social Science
Exploring how to treat LLMs as autonomous agents in specific scenarios lead to diverse and promising applications in social science [30, 34, 33, 10]. For instance, Park et al. established an LLM-powered agent framework to simulate human behavior in an interactive scenario, demonstrating the potential of LLMs to model complex social interactions and decision-making processes. Moreover, the application of LLM agents in economic research has been explored, providing new insights into financial markets and economies [11, 17]. Extending beyond the realm of social sciences, Mao et al. adeptly utilized LLMs to generate driving trajectories in motion planning tasks. In the field of natural sciences, Williams et al. integrated LLMs with epidemic models to simulate the spread of diseases. These varied applications highlight the versatility and potential of LLMs to understand and model various real-world dynamics.
@@ 3 Methodology
We consider the generation of individual daily activity trajectories, each representing an individual’s activities for the whole day. In addition, we focus on the urban context, where the activity trajectory of each individual is represented as a time-ordered sequence of location choices (e.g., POIs) [20]. This sequence is represented by $\{(l_{0},t_{0}),(l_{1},t_{1}),\ldots,(l_{n},t_{n})\}$, where each $(l_{i},t_{i})$ denotes the individual’s location $l_{i}$ at time $t_{i}$.
Figure 2: LLMob, the proposed LLM agent framework for personal Mobility generation.
By modeling individuals within an urban environment as LLM agents, we present LLMob, an LLM Agent Framework for Personal Mobility Generation, as illustrated in Figure 2. LLMob is based on the assumption that an individual’s activities are primarily influenced by two principal factors: habitual activity patterns and current motivations. Habitual activity patterns, representing typical movement behaviors and preferences that indicate regular travel and location choices, are recognized as crucial information for inferring daily activities [7, 29]. On the other hand, motivations relate to dynamic and situational elements that sway an individual’s choices at any particular moment, such as immediate needs or external circumstances during a specific period. This consideration is vital for capturing and forecasting short-term shifts in mobility patterns [1, 38]. Moreover, by formulating prompts that assume specific events of concern, this framework allows us to observe the LLM agent’s responses in a variety of situations.
To construct a pipeline for activity trajectory generation, we design an LLM agent with action, memory, and planning [31, 30]. Action specifies how an agent interacts with the environment and makes decisions. In LLMob, the environment contains the information collected from real-world data, and the agent act by generating trajectories. Memory includes past actions that need to be prompted to the LLM to invoke the next action. In LLMob, memory refers to the patterns and motivations output by the agent. Planning formulates or refines a plan over past actions to handle complex tasks, with additional information optionally incorporated as feedback. In LLMob, we use planning to identify patterns and motivations, thereby handling the complex task of trajectory generation. Plan formulation, selection, reflection, and refinement [14] are employed in succession, and the agent keeps updating the action plan based on its observation [36]: The agent first formulates a set of activity plans by extracting candidate patterns from historical trajectories in the database. Then, the agent performs self-reflection through a self-consistency evaluation to pick the best pattern from the candidate patterns. With historical trajectories further retrieved from the database, the agent refines the identified pattern to a summarized motivation of daily activity, which is then jointly used with the identified pattern for trajectory generation. Besides the above agentic components, we also specify personas to the agent to simulate the diversity of real-world individuals [28].
3.1 Activity Pattern Identification
Phase 1 of LLMob focuses on identifying activity patterns from historical data. To effectively leverage the extracted activity patterns as essential prior knowledge for the generation of daily activities, we introduce the following two steps.
@@ 3.1.1 Pattern Extraction from Semantics and Historical Data
This step derives activity patterns based on activity trajectory data (e.g. individual check-in data). As illustrated in the left panel of Figure 2, this scheme consists of the following aspects: For each person, we start by specifying a candidate personas to the LLM agent, providing the inspiring foundation for subsequent activity pattern generation. This approach also encourages the diversity of the generated activity patterns, as each candidate persona acts as a unique prior for the generation process (e.g., the significance of user clustering from activity trajectory data in producing meaningful distinctions has been demonstrated [23]). Meanwhile, we perform data preprocessing to extract key information from the extensive historical data. This involves identifying usual commuting distances, pinpointing typical start and end times and locations of daily trips, and concluding the most frequently visited locations of the person. It is important to note that these pieces of information are widely recognized as critical features in mobility analysis [15]. Following this, both semantic elements with historical data are combined in the prompts, requiring the LLM agent to summarize the activity patterns for this person. By doing this, we set up a streamline to effectively bridge the gap between semantic persona characteristics and concrete historical activity trajectory data, which allows for a more personalized and interpretable representation of individual activities at one day. For each person, a set of $C$ ($C=10$) candidate patterns, denoted as $\mathcal{CP}$, are generated according to the historical data and $C$ candidate personas, respectively.
@@ 3.1.2 Pattern Evaluation with Self-Consistency
This step involves assessing the consistency of the candidate patterns to identify the most plausible one. We implement a scoring mechanism to gauge the alignment of candidate patterns with historical data. To achieve this objective, we define a scoring function for each candidate pattern $cp$ in the set $\mathcal{CP}$. This function evaluates $cp$ against two distinct sets of activity trajectories: the specific activity trajectories $\mathcal{T}_{i}$ of a targeted resident $i$ and the sampled activity trajectories from other residents $\mathcal{T}_{\sim i}$:
$score_{cp}=\sum_{t\in\mathcal{T}_{i}}r_{t}-\sum_{t^{\prime}\in\mathcal{T}_{\sim i}}r_{t^{\prime}}$ (1)
where we design an evaluation prompt to ask the LLM to generate rating scores $r_{t}$ and $r_{t^{\prime}}$. Specifically, the LLM agent is prompted to assess the degree of preference for a given trajectory based on the candidate pattern. Ideally, the LLM agent should assign a higher $r_{t}$ for data from the targeted resident and a lower $r_{t^{\prime}}$ for data from other residents. This scheme requires self-consistency in the pattern: the activity pattern derived from the activity trajectory data should be consistent with the data during the evaluation. We provide the pseudo-code of the algorithm for Phase 1 of LLMob in Appendix A.